OpenStudy (anonymous):

the midpoint of AB is (3,-5) A =(9,4) What are the coordinates for B. I know there is a trick to figuring this out by adding/subtracting the individual coordinates by each other. Any help?

OpenStudy (anonymous):

@galacticwavesXX

OpenStudy (anonymous):

ok you know the midpoint formula is \[M=(\frac{ x1+x2 }{ 2 },\frac{ y1+y2 }{ 2 })\]

OpenStudy (anonymous):

\[M=AB=(3,-5)\] Mx=3 & My=-5 so, \[Mx=\frac{ x1+x2 }{ 2 }, My=\frac{ y1+y2 }{ 2 }\]

OpenStudy (anonymous):

you understand what i've done so far?

OpenStudy (anonymous):

Yes you separated the midpoint coordinates for x and y

OpenStudy (anonymous):

yup now we know what x1 & y1 are and also what Mx & My is. Now all we have to do is plug them in to the designated equation and solve for x2 & y2

OpenStudy (anonymous):

Tell me what you get for x2 & y2 when you're finished solving for the problem

OpenStudy (anonymous):

Im lost I got -5=(9+y2)/2 for My

OpenStudy (anonymous):

no it should be \[My=-5=\frac{ 4+y2 }{ 2 }\] you are solving for y2, so you know you have to set the equation to y2

OpenStudy (anonymous):

x1=9 & y1=4. They should only be used in their respective formulas.

OpenStudy (anonymous):

Still confused sorry

OpenStudy (anonymous):

\[Mx=\frac{ x1+x2 }{ 2 }\rightarrow3=\frac{ 9+x2 }{ 2 }\] set everything to x2 so the equation will be, \[(3*2)=9+x2\rightarrow6=9+x2\rightarrow-3=x2\] you can now see x2= -3 \[My=\frac{ y1+y2 }{ 2 }\rightarrow-5=\frac{ 4+y2 }{ 2 }\] solved for y2, the same way you solved for x2. \[(-5*2)=4+y2\rightarrow-10=4+y2\rightarrow-6=y2\] you can see that y2 = -6 so, \[B=(-3,-6)\]

OpenStudy (anonymous):

\[M=(Mx,My)=AB\] just to help

OpenStudy (anonymous):

woahhhhhhhhh it should be B=(-3,-14)

OpenStudy (anonymous):

lol arithmetic error

OpenStudy (anonymous):

Yea I saw, -5-9=13

OpenStudy (anonymous):

-13* Thanks

OpenStudy (anonymous):

lol -5-9=-14 but yea sign always gets me, most irritating mistake to make in math